OpenStudy (anonymous):

How many two digit numbers are divisible by 2 and 5 ( Not apart of homework, just going over Hunter Sample Entrance Exam with different numbers)? -Please provide with a formula.

OpenStudy (anonymous):

these are A.P.series 10+12+.....+98 find n1 ,the number of numbers divisible by 2 10+15+...+95 find n2 ,the number of numbers divisible by 5 10+20+...+90 find n3 ,the number of numbers divisible by 10 then required number =n1+n2-n3

OpenStudy (anonymous):

So, it's 54.

OpenStudy (anonymous):

@surjithayer

OpenStudy (radar):

No, think about it. two digit numbers will range from 10 to 99, Of those you only count those that are divisible by 5 AND 2. The number divided by 5 is only 18 and only half of them are divided by 2.

OpenStudy (mathstudent55):

If a number is divisible by 2 and 5, it must have at least 2 and 5 as factors. The smallest number, then, is 2 * 5 = 10 Now find every multiple of 10 that is a 2-digit number. 10 * 1 10 * 2 10 * 3 etc.

OpenStudy (anonymous):

There are 90 two digit numbers.

OpenStudy (anonymous):

Wouldn't 90/10= 9?, so then there are 9 multiples of 10 within 90.

OpenStudy (radar):

Good reasoning, there are 9 numbers from 10 to 99 that can be divided (evenly) by both 5 and 2.

OpenStudy (anonymous):

54 is correct answer.

OpenStudy (anonymous):

45+18-9=63-9=54

OpenStudy (radar):

I believe the answer of 54 is looking at the problem as if it had asked " ...divisible by 2 OR 5 rather than 2 AND 5. The "and" function would require each counted two digit number to be divisible by both 2 and 5 and for the life of me I can only count 9 which are 10, 20, 30, 40, 50, 60, 70, 80, and 90. existing from 10 to 99. Any others will not be divisible by both 2 and 5.

OpenStudy (mathstudent55):

I agree with @radar

OpenStudy (anonymous):

i also agree with you when it is 2 and 5 means divisible by 10